OpenStudy (secret-ninja):

Medals. Can someone walk me through how to do this? I really need an expert!!! A system of equations is shown below: 8x + 5y = 9 3x + 2y = 4 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (secret-ninja):

@amistre64 @thadyoung @undeadknight26

undeadknight26 (undeadknight26):

I suck at these things ._.

OpenStudy (secret-ninja):

Darn .:[

OpenStudy (isaiah.feynman):

Hmmm if you know the solutions to this system you can create ANY system of equations. But we have to follow the rules here.

OpenStudy (secret-ninja):

What are the rules?

OpenStudy (secret-ninja):

"replacing one equation with the sum of that equation and a multiple of the other" you mean?

OpenStudy (isaiah.feynman):

Yes. I don't understand that statement.

OpenStudy (secret-ninja):

Me neither. One of my relatives is verrry VERY great at math and even she didn't know, so i really feel hopeless with this.

OpenStudy (secret-ninja):

I guess I'll just close this and fail. :/ I never learned this. v-v

OpenStudy (isaiah.feynman):

@Secret-Ninja

OpenStudy (secret-ninja):

Yes?

OpenStudy (isaiah.feynman):

I'm pondering.

OpenStudy (secret-ninja):

Okay.

OpenStudy (secret-ninja):

https://answers.yahoo.com/question/index?qid=20140719095908AAMsup4 I found this but its all messed up and I can't make any sense of it. :/

OpenStudy (isaiah.feynman):

Gotcha!

OpenStudy (isaiah.feynman):

Thats exactly what I was thinking!

OpenStudy (secret-ninja):

?

OpenStudy (mathmath333):

example multiple of \(\Large 7=n(7)\\ where\\n=1,2,3.............\) just exactly the multiple of any equation would be \(\Large n(ax+by)=n(c)\)

OpenStudy (secret-ninja):

Sorry... I still don't get it. o.o

OpenStudy (secret-ninja):

Part (A) is easy if you don't need the answer to be useful. Just follow the steps. Replace the first equation with the sum of it and a multiple of the other equation. Suppose the multiple is -2 (I want to make the numbers smaller, not larger, but either way works.) Then: 8x + 5y = 9 .... original equation -6x - 4y = -8 .... -2 times each term in the 2nd equation -------------------- 2x + y = 1 .... "sum" of the above, term-by-term on both sides That leaves you with a new system: 2x + y = 1 .... first equation replaced 3x + 2y = 4 .... 2nd equation unchanged

OpenStudy (secret-ninja):

Thats what the other person said...

OpenStudy (mathmath333):

example \(\Large x+2y=3---A\) \(\Large 2(x+2y)=2(3)\\ \Large 2x+4y=6-----B\) if we try to put \(\Large x=2\) say then from \(\Large A\) and \(\Large B\) both u will get \(\Large y=\dfrac{1}{2}\)

OpenStudy (secret-ninja):

@satellite73 Please help

OpenStudy (secret-ninja):

I know you'll proably totally hate me for this... but I still don't get it. sowwy v-v

OpenStudy (secret-ninja):

I'm just going to write a bunch or rubbish.... Even if I get this wrong I'll get a B on the test.

OpenStudy (secret-ninja):

*of

OpenStudy (isaiah.feynman):

Okay..